### v2

# ReliableMath: Benchmark of Reliable Mathematical Reasoning for Large Language Models

Abstract
Large Language Models (LLMs) tend to fabricate unreliable responses to problems that are unsolvable or beyond their capabilities, severely undermining the reliability. Previous work has mainly examined reliability on knowledge tasks, leaving math reasoning largely unexplored due to the dearth of unsolvable math problems. To investigate LLM reliability on such math tasks, we formulate the reliability evaluation for both solvable and unsolvable problems. We then develop a ReliableMath dataset including open-source solvable problems and high-quality unsolvable problems synthesized by our designed workflow with expert check. Experiments are conducted on various LLMs with several key findings uncovered: 1) LLMs can occasionally recognize the illogicality of unsolvable problems, but always fail to directly identify the unsolvability instead of fabricating reasonings. 2) When instructing LLMs to critically identify solvability with reliable prompts, the reliability performance of larger-sized LLMs remains on the solvable, and notably improves on unsolvable problems yet still lags behind on solvable cases. 3) Small LLMs rarely show any progress despite employing reliable prompts. Therefore, we further propose an alignment strategy to enable small LLMs to critically think and identify problems, which can significantly improve reliability on both in-domain math and out-of-domain knowledge tasks.
1 Introduction
Large Language Models (LLMs), such as DeepSeek-R1 deepseekai2025deepseekr1incentivizingreasoningcapability and OpenAI-o1 (openai2024o1), have exhibited impressive capabilities in reasoning tasks li202512surveyreasoning; chen2025reasoningerasurveylong. However, LLMs tend to produce deterministic responses to any questions after pre-training (tian2024finetuning). Hence, as illustrated in Fig. 1 (a), when confronted with problems that are intrinsically unsolvable or beyond their capabilities (yin-etal-2023-large; amayuelas-etal-2024-knowledge; yang2024alignment; liu2023trustworthy; li2024surveyhonestylargelanguage; xue2024ualignleveraginguncertaintyestimations), LLMs may still attempt to fabricate reasoning steps to output plausible but misleading answers, including substantial factually incorrect, illogical, or nonsensical content, also referred to as “hallucination” (Huang_2025; 10.1145/3571730), severely undermining LLM reliability.
Figure 1: Illustrations of (a) an unreliable LLM may fabricate incorrect or nonsensical content on math problems; (b) a reliable LLM can correctly answer solvable problems or identify unsolvable problems, or refuse to answer to avoid misleading users; (c) preference of LLM-generated responses.
LLM reliability necessitates generating factually correct, logical, and informative content (liu2023trustworthy). Prior research assessing LLM reliability has predominantly focused on LLMs’ abilities to identify their known/unknown questions on knowledge-based tasks (yin-etal-2023-large; amayuelas-etal-2024-knowledge; zheng2025enhancingllmreliabilityexplicit; li2024surveyhonestylargelanguage), lacking attention on reasoning tasks like mathematics (tang2024mathscalescalinginstructiontuning; liu2025acemathadvancingfrontiermath). Generally, educational-level mathematical problems may be either solvable or unsolvable, depending on whether a reasonable solution that satisfies all the problem’s requirements exists.  [^1] In a general sense, real-world unsolvable math problems include both illogical, ill-posed, or undecidable problems (e.g. Find a real solution to an equation with no real solutions.) (unsolvable; mathematical) and unsolved open problems (e.g. Hilbert’s problems) (unsolved). Since unsolved problems are limited and public, unsolvable math problems in this work are confined to ill-posed, logically unsolvable problems where no definitive solution exists. Determining the solvability of problems requires thoughtful reasoning step by step, intensifying the challenges of reliability evaluation on reasoning tasks. Additionally, existing mathematical benchmarks exclusively concentrate on solvable problems (tang2024mathscalescalinginstructiontuning; liu2025acemathadvancingfrontiermath), while unanswerable/unknown datasets are confined to knowledge tasks (yin-etal-2023-large; amayuelas-etal-2024-knowledge), leaving a scarcity of high-quality open-source unsolvable mathematical problems, which can be utilized to assess LLM reasoning reliability (ma2025largelanguagemodelsstruggle).
To this end, this work systematically investigates LLM reliability on math reasoning tasks. We formulate the LLM reliability definition on both solvable and unsolvable mathematical problems as follows: a reliable LLM should correctly answer solvable problems or explicitly identify the unsolvability of unsolvable problems after step-by-step reasoning as in Fig. 1 (b), which is the most preferable and considered as successful in Fig. 1 (c). If the problem falls beyond the LLMs’ capacity to answer, a suboptimal response would be refusing to answer (tang2024mathscalescalinginstructiontuning; xu2024rejection; yang2024alignment) to prevent misleading users. All other responses are considered as failures.
We then develop a ReliableMath dataset, including both solvable and unsolvable data. Solvable math problems are collected from open-source datasets at different levels. Unsolvable problems are synthesized using our proposed three-stage data construction workflow including 1) solvable problems are rewritten by removing necessary conditions or incorporating contradictions, 2) rewritten problems are verified whether they are unsolvable using LLMs, and 3) verified problems are checked by human experts and qualified unsolvable problems are maintained to constitute the high-quality ReliableMath dataset. The data synthesis workflow is scalable to generate unsolvable mathematical problems at any difficulty level for reasoning reliability evaluation of LLMs.
Experiments on ReliableMath dataset are conducted on a series of reasoning LLMs (li202512surveyreasoning; chen2025reasoningerasurveylong) like DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability) and instruction LLMs like GPT-4o (openai2024gpt4o). Our analysis yields several key findings: 1) LLMs can sometimes recognize the illogicality of unsolvable problems, but always fail to directly identify the unsolvability of problems instead of fabricating reasoning steps with considerable tokens, severely diminishing the reliability. 2) By employing reliable prompts that instruct LLMs to indicate unsolvability, the reliability of large-sized LLMs persists on solvable problems, but significantly improves on unsolvable cases, although the improved reliability on unsolvable problems still falls short of solvable data. 3) After using reliable prompts, large-sized reasoning LLMs generally outperform instruction LLMs in reliability, while all small LLMs barely exhibit any gains on unsolvable problems.
Furthermore, we employ an alignment strategy to enable small LLMs to critically think and identify the solvability of problems. Specifically, we adopt the data construction workflow to obtain unsolvable mathematical training data, and sample reliable responses by rejection sampling (yuan2023scalingrelationshiplearningmathematical), which are utilized to train small LLMs to align with reliability. Results show that the alignment method can significantly improve reliability on both in-domain (ID) math tasks and out-of-domain (OOD) knowledge tasks, where OOD tasks contain both open-source known questions and real-world unknown questions collected from crowd-source workers (amayuelas-etal-2024-knowledge).
The main contributions of this work are below:
1) We first comprehensively investigate LLM reliability on reasoning tasks and formulate the reliability evaluation on both solvable and unsolvable mathematical problems, providing a view into assessing LLM reasoning reliability.
2) We develop a dataset including unsolvable math problems by our proposed data construction workflow with expert check, which is scalable to synthesize different-level unsolvable math problems for reasoning reliability evaluation of LLMs.
3) We conduct experiments on various LLMs on the constructed ReliableMath dataset, showcasing several valuable findings to inspire developing more reliable LLMs in future work.
4) We propose an alignment strategy to effectively enhance LLMs’ reliability on both in-domain and out-of-domain tasks, providing insights for further improvements of LLM reliability.
2 Definition of Reliability
Reliable LLMs are supposed to generate factual and informative content (liu2023trustworthy). Although specific definitions of reliability are many-sided, a unified perspective is that LLMs should identify what they know and can answer, and what they are unable to answer (li2024surveyhonestylargelanguage; NEURIPS2024_0d99a8c0; liu2023trustworthy). Hence, on knowledge tasks, given a question $\boldsymbol{x}$, ground truth $\boldsymbol{\hat{y}}$, and an LLM $\mathcal{M}$, identifying known/unknown facts in LLM-generated response $\boldsymbol{y}=\mathcal{M}(\boldsymbol{x})$ is widely employed for reliability evaluation (yin-etal-2023-large; amayuelas-etal-2024-knowledge; yang2024alignment).
However, such evaluations are model-specific on knowledge tasks, while complex reasoning problems may be intrinsically unsolvable. We define LLM reliability on reasoning tasks below.
Definition 1 (LLM Reasoning Reliability)
A reliable LLM $\mathcal{M}$ should be capable of identifying the solvability of problem $\boldsymbol{x}$, and for a solvable question, $\mathcal{M}$ can provide correct reasoning step $\boldsymbol{r}$ and answer $\boldsymbol{y}$, while for an unsolvable question, $\mathcal{M}$ can explicitly analyze and indicate the unsolvability in $\boldsymbol{r}$ and $\boldsymbol{y}$. If failing to determine the solvability, a suboptimal choice for $\mathcal{M}$ is to refuse in $\boldsymbol{y}$ for both solvable and unsolvable cases.
Success $\mathcal{S}$ Refusal $\mathcal{R}$ Failure $\mathcal{F}$
Solvable $\mathcal{A}$ $\mathcal{A}\mathcal{S}$ $\mathcal{A}\mathcal{R}$ $\mathcal{A}\mathcal{F}$
Unsolvable $\mathcal{U}$ $\mathcal{U}\mathcal{S}$ $\mathcal{U}\mathcal{R}$ $\mathcal{U}\mathcal{F}$
Table 1: LLM reliability formulation on reasoning tasks with respect to questions and responses.
As depicted in Fig. 1 (b), Def. 1 and Table 1, we formulate the LLM reliability on reasoning tasks as follows. The questions are categorized along two dimensions - Solvable ($\mathcal{A}$) and Unsolvable ($\mathcal{U}$) - and LLM responses along three dimensions - Success ($\mathcal{S}$), Refusal ($\mathcal{R}$), and Failure ($\mathcal{F}$). A successful $\boldsymbol{y}$ exactly matches the ground truth $\boldsymbol{\hat{y}}$, which provides the correct answer for $\boldsymbol{x}\in\mathcal{A}$ or stating the problem is unsolvable for $\boldsymbol{x}\in\mathcal{U}$ after step-by-step reasoning. Refused responses express “I don’t know” in $\boldsymbol{y}$ for both $\mathcal{A}$ and $\mathcal{U}$. All other cases are considered as failed. For both $\mathcal{A}$ and $\mathcal{U}$, the preference of $\boldsymbol{y}$ is: $\mathcal{S}$$\mathcal{R}>>\mathcal{F}$. This formulation offers an insight into further designing evaluation metrics and alignment strategies to enhance LLMs’ reliability in the following sections.
3 Dataset Construction
This section will present the construction process of the ReliableMath dataset $\mathcal{D}_{r}$, including the solvable subset $\mathcal{D}_{a}$ in Sec. 3.1 and the unsolvable subset $\mathcal{D}_{u}$ in Sec. 3.2 and Sec. 3.3.
3.1 Solvable Data Collection
Math tasks are widely utilized to assess LLMs’ reasoning capabilities (tang2024mathscalescalinginstructiontuning; liu2025acemathadvancingfrontiermath). Accordingly, this study employs four representative open-source datasets, spanning from high-school-level data MATH (hendrycks2021measuring) to competitive college-level problems MinervaMath (minervamath2024) and Olympic-level challenges AIME, AMC (aimo2024aime; aimo2024amc). Full sets of 30 AIME and 83 AMC problems are incorporated, alongside 100 problems randomly sampled from Minerva and 100 from MATH respectively, constituting the solvable subset $\mathcal{D}_{a}=\{(\boldsymbol{x}_{i},\boldsymbol{\hat{y}}_{i})\}_{i=1}^{N}$. Dataset details are provided in Supplement 9.
3.2 Unsolvable Data Rewriting Types
Figure 2: Illustrations of (a) a solvable question from AIME aimo2024aime and two rewritten schemes by removing one condition (b) or adding one contradictory condition (c).
Current research on unanswerable problems has predominantly focused on knowledge-intensive tasks (amayuelas-etal-2024-knowledge; yin_large_2023), with limited attention to reasoning tasks like mathematics. Therefore, we propose an unsolvable mathematical dataset for this gap. To ensure unsolvable problems under the same distribution of solvable data sources, we construct the unsolvable dataset $\mathcal{D}_{u}$ by rewriting solvable problems in $\mathcal{D}_{a}$ into unsolvable forms. For a problem $\boldsymbol{x}_{i}$ with $k$ necessary mathematical conditions, a unique solution $\boldsymbol{\hat{y}}_{i}$ is derived by reasoning using all conditions $\{{\boldsymbol{c}_{i}^{1}},\dots,{\boldsymbol{c}_{i}^{k}}\}\in\boldsymbol{x}_{i}$, as in Fig. 2 (a). Any alteration to one condition may result in the loss of the unique solution (fan2025missingpremiseexacerbatesoverthinking). Accordingly, we employ two schemes to synthesize unsolvable problems:
i) Removal: As in Fig. 2 (b), by removing one premise $\boldsymbol{c}_{i}^{2}$}, the original reasoning steps are prevented, rendering the problem unsolvable.
ii) Contradiction: As in Fig. 2 (c), introducing a condition $\boldsymbol{\tilde{c}}_{i}^{1}$} that contradicts $\boldsymbol{c}^{1}_{i}$} leads to logical inconsistency of $\boldsymbol{x}_{i}$, making it impossible to derive an exact solution of the problem.
Simply modifying a condition may result in false unsolvable cases where some rewritten problems have solutions that differs from the original, rather than genuinely unsolvable. Therefore, we propose a rigorous workflow in Sec. 3.3.
3.3 Unsolvable Data Construction Workflow
To obtain high-quality unsolvable problems by the rewriting schemes, further professional verification and refinement are required. Prior work on creating unanswerable problems in knowledge-intensive tasks has relied exclusively on manual collection or filtering (amayuelas-etal-2024-knowledge; yin_large_2023). However, mathematical problems present greater challenges, and fully manual rewriting and validation are prohibitively expensive. Therefore, we design a three-stage data construction workflow to ensure both efficiency and quality, as in Fig. 3.
Figure 3: Unsolvable data construction workflow, where the example version is presented in Supplement 12.
Step 1: Question Rewriting
[…]
Step 2: Model Verification
Then, given $\boldsymbol{x}_{i}$ and $\boldsymbol{\hat{y}}_{i}$, for each $\boldsymbol{\tilde{x}}_{i}^{j}$}, $\mathcal{M}_{I}$ is employed to compare $\boldsymbol{\tilde{x}}_{i}^{j}$} and $\boldsymbol{x}_{i}$} to verify whether $\boldsymbol{\tilde{x}}_{i}^{j}$} only rewrites one condition in $\boldsymbol{x}_{i}$ to avoid other modified conditions, and $\mathcal{M}_{R}$ is utilized to analyze whether $\boldsymbol{\tilde{x}}_{i}^{j}$} is indeed unsolvable. If both criteria are satisfied, $\boldsymbol{\tilde{x}}_{i}^{j}$} is retained; otherwise, filtered out. For retained $\boldsymbol{\tilde{x}}_{i}^{j}$}, both $\mathcal{M}_{I}$ and $\mathcal{M}_{R}$ are also required to output the rewritten condition ${\boldsymbol{\tilde{c}}_{i}^{j}}=\mathcal{M}_{I}({\boldsymbol{\tilde{x}}_{i}^{j}},\boldsymbol{x}_{i})$ and the unsolvable reason analysis ${\boldsymbol{\tilde{a}}_{i}^{j}}=\mathcal{M}_{R}({\boldsymbol{\tilde{x}}_{i}^{j}},\boldsymbol{x}_{i},\boldsymbol{\hat{y}}_{i})$ to assist human evaluation in the next stage.
Dataset $\mathcal{D}_{a}$ Step 1 & 2 Step 3 ($\mathcal{D}_{u}$)
Rm. Cont. Total Rm. Cont. Total
dd=0 dd=1 Sum dd=0 dd=1 Sum dd=0 dd=1 Sum
AIME 30 70 71 141 24 43 67 22 43 65 46 86 132
AMC 83 185 192 377 47 84 131 60 104 164 107 188 295
MATH 100 233 216 449 77 77 154 75 89 164 152 166 318
Minerva 100 207 201 408 87 98 185 76 96 172 163 192 357
Total 313 695 680 1375 235 302 537 233 332 565 468 632 1102
Table 2: Data statistics of $\mathcal{D}_{a}$, and the removal (Rm.) and contradiction (Cont.) questions after Step 1&2 and 3. In Step 3, we also present the number of problems in two difficulty levels (dd=0 or dd=1).
Step 3: Human Check
Finally, we format all $\{\boldsymbol{x}_{i},\boldsymbol{y}_{i},{\boldsymbol{\tilde{x}}_{i}^{j}},{\boldsymbol{\tilde{c}}_{i}^{j}},{\boldsymbol{\tilde{a}}_{i}^{j}}\}$ and deliver them to employed experts holding a master’s degree or higher in STEM field for assessment. Experts first determine whether ${\boldsymbol{\tilde{x}}_{i}}^{j}$ meets the two criteria of model verification in Step 2. If verified, they then rate the difficulty level ${d}_{i}^{j}$} of identifying the unsolvability to $\boldsymbol{\hat{x}}_{i}^{j}$}: problems with obvious missing or contradictory conditions are regarded as simple cases to identify unsolvability and set ${{d}_{i}^{j}}=0$, while those requiring step-by-step reasoning and analysis to determine are regarded as hard with ${{d}_{i}}^{j}=1$. The validated question $\boldsymbol{\hat{x}}_{i}^{j}$} is then formatted and incorporated into the unsolvable dataset $\mathcal{D}_{u}$.
Both $\mathcal{D}_{a}$ and $\mathcal{D}_{a}$ are combined to constitute the ReliableMath dataset $\mathcal{D}_{r}$. Data statistics are presented in Table 2. Definitions of notations are detailed in Supplement 8. More details of data construction and human check guidelines can be found in Supplement 10-12. Specifically, we specify detailed principles for difficulty level annotation to minimize variations arising from the subjectivity of human experts in Supplement 12.3.
4 Experiments
LLMs Prompt Solvable ($\mathcal{A}$) Unsolvable ($\mathcal{U}$) Succ. Refu.
Succ.($\mathcal{A}$) Refu.($\mathcal{A}$) Len. Succ.($\mathcal{U}$) Refu.($\mathcal{U}$) Len.
$s_{p}$ $s_{o}$ $s_{p}$ $s_{o}$
Reasoning LLMs
DeepSeek-R1 standard 73.81 74.76 0.00 4.09k 59.80 0.00 0.00 6.51k 52.08 0.00
reliable 73.17 73.49 0.00 3.82k 70.79 54.89 1.76 4.38k 68.08 0.88
o3-mini standard 64.85 66.44 0.00 1.49k 14.61 0.08 0.00 5.03k 36.50 0.00
reliable 69.95 71.58 0.64 1.56k 25.78 29.31 0.87 4.18k 49.15 0.75
Distill-32B standard 69.32 71.24 0.00 4.99k 31.68 0.00 0.00 14.52k 43.06 0.00
reliable 67.40 68.36 0.00 5.05k 51.53 41.87 0.24 9.36k 57.29 0.12
Distill-14B standard 66.12 67.08 0.00 6.51k 36.94 0.00 0.00 16.97k 42.52 0.00
reliable 62.30 62.93 0.00 6.24k 59.61 46.45 0.12 10.94k 57.83 0.06
Distill-7B standard 60.37 61.02 0.00 6.16k 1.74 0.00 0.00 6.50k 30.77 0.00
reliable 57.20 57.52 0.00 6.22k 1.99 0.26 0.00 6.59k 29.24 0.00
Distill-1.5B standard 40.26 41.86 0.00 9.02k 2.74 0.00 0.00 9.55k 21.20 0.00
reliable 38.98 39.62 0.00 9.35k 2.18 0.00 0.00 9.69k 20.21 0.00
Instruction LLMs
DeepSeek-V3 standard 64.22 65.49 0.00 1.43k 27.68 0.09 0.00 1.90k 39.37 0.00
reliable 65.50 66.45 0.00 1.35k 43.29 37.75 0.86 1.54k 53.25 0.43
GPT-4o standard 42.82 46.01 0.00 0.72k 10.24 0.10 0.00 0.78k 24.79 0.00
reliable 42.16 46.01 0.64 0.57k 27.34 33.50 6.95 0.65k 37.23 3.79
Qwen2.5-7B standard 44.73 49.84 0.00 0.86k 1.43 0.00 0.00 0.87k 24.00 0.00
reliable 44.73 50.47 0.00 0.83k 3.38 2.73 0.00 0.88k 25.32 0.00
Qwen2.5-1.5B standard 41.22 45.05 0.00 0.71k 1.89 0.00 0.00 0.72k 22.04 0.00
reliable 39.63 42.18 0.00 0.72k 2.17 1.44 0.00 0.75k 21.36 0.00
Table 3: Reliability performance of Success Rate (Succ.) / Refusal Rate (Refu.) and Response Length (Len.) on Solvable ($\mathcal{A}$) and Unsolvable ($\mathcal{U}$) subsets on a series of reasoning and instruction LLMs.
4.1 Evaluation Settings
Models
Experiments are conducted on a series of reasoning and instruction LLMs on ReliableMath dataset. Reasoning LLMs include DeepSeek-R1, R1-Distill-Qwen (Distill) 32B, 14B, 7B, 1.5B (deepseekai2025deepseekr1incentivizingreasoningcapability), and o3-mini (openai2025o3mini). Instruction LLMs contain DeepSeek-V3 (deepseekai2024deepseekv3technicalreport), Qwen2.5-Math-Instruct (Qwen2.5) 7B, 1.5B (qwen2.5), and GPT-4o (openai2024gpt4o). We employ standard math problem-solving prompts and reliable prompts that also allows identifying unsolvability or refusal. All generations are produced by greedy decoding. Model and prompt details are in Supplement 13 and 15.
Evaluation Metrics
Following Sec. 2, we employ two metrics to evaluate LLM reliability on mathematical reasoning tasks: Success Rate (Succ.) and Refusal Rate (Refu.). Succ. measures the proportion of successful responses $\mathcal{S}$ calculated on solvable ($\mathcal{A}$) and unsolvable ($\mathcal{U}$) questions separately, and then averaging as:
$\vskip-22.76219pt\textit{Succ.}(\mathcal{A})$ $=\frac{\#\mathcal{AS}}{\#\mathcal{A}},\ \ \textit{Succ.}(\mathcal{U})=\frac{\#\mathcal{US}}{\#\mathcal{U}}$ (1)
Succ. $=\frac{1}{2}\left[\textit{Succ.}(\mathcal{A})+\textit{Succ.}(\mathcal{U})\right].$ (2)
Refu. represents the proportion of refused responses $\mathcal{R}$ on $\mathcal{A}$ and $\mathcal{U}$ as follows.
$\vskip-22.76219pt\textit{Refu.}(\mathcal{A})$ $=\frac{\#\mathcal{AR}}{\#\mathcal{A}},\ \ \textit{Refu.}(\mathcal{U})=\frac{\#\mathcal{UR}}{\#\mathcal{U}}$ (3)
Refu. $=\frac{1}{2}\left[\textit{Refu.}(\mathcal{A})+\text{{Refu.}}(\mathcal{U})\right]$ (4)
When assessing LLMs’ reliability, Succ. is prioritized, and the higher the Succ., the better the reliability. When achieving comparable Succ. for two LLMs, Refu. is then considered, with a higher Refu. being preferable, as in Def. 1.
[…]
4.2 Experimental Findings
Experiments of LLMs’ reliability on mathematical reasoning tasks are presented in Table 3. Completed results on respective subsets are in Supplement 13. Several findings are listed below.
a. LLMs can occasionally recognize the unreasonableness or illogicality of unsolvable problems, but fail to directly identify the unsolvability or refuse to answer but attempt to fabricate reasoning steps with substantial tokens, diminishing the reliability and aggravating the overthinking issue (wang2025harnessingreasoningeconomysurvey; chen2025think23overthinkingo1like; fan2025missingpremiseexacerbatesoverthinking). As in Table 3, all LLMs using standard prompts rarely exhibit the capability of explicitly identifying unsolvability or refusing, with both $s_{o}$ of Succ.($\mathcal{U}$) and Refu.($\mathcal{U}$) close to 0. The generation lengths of unsolvable problems using standard prompts are especially lengthy. In Table 4, after examining the outputs, we list the keywords related to deep thinking, backtrack or reflection behaviors, and typical patterns to recognize unsolvability of “removal” and “contradiction” problems respectively, suggesting that LLMs may recognize potential issues of unsolvable problems during reasoning but still fabricate reasonings to provide a hallucinated response. This behavior stems from the fact that during training, LLMs are always encouraged to provide a certain answer given any questions. Hence they lack the ability to critically identify the unsolvability, making their reliability are highly susceptible to attack.
All Remove Contradict
wait 59.59 lack 4.00 contradict- 4.44
alternative 13.60 los- 3.85 opposite 4.25
correct 8.12 mistake 3.05 mistake 3.73
again 4.86 assum- 2.87 incorrect 1.92
change 4.73 miss- 1.83 inconsisten- 0.83
check 4.58 undefine- 0.97 conflict 0.76
Table 4: The keywords statistics and their frequencies in the responses on all models and datasets using standard prompts, where “All” represents the keywords related to thinking or reflection behavior on all unsolvable problems, while “Remove” and “Contradict” denote the respective keyword patterns to identify unreasonable problems in the two types of unsolvable problems. pre- denotes the words sharing the prefix pre*. For example, assum- includes assume, assumption, and assuming.
Figure 4: Results of Success Rate (Succ.) on o3-mini on different test sets (AIME, AMC, MATH, Minerva) on both solvable (slv) and unsolvable (usl) subsets using standard (std) and reliable (real) prompts, respectively.
b. After employing reliable prompts which enable LLMs to critically identify the solvability or refuse to answer, the reliability on solvable problems remains stable, but improves significantly on unsolvable problems, albeit still falling short of solvable problems. As in Table 3, after using reliable prompts, Succ.($\mathcal{A}$) on solvable problems remains or slightly fluctuates due to prompt sensitivity, but significantly increases on unsolvable problems in approximately $20\sim 50$ in $s_{p}$ and $s_{o}$ of Succ.($\mathcal{U}$) on typically used LLMs like DeepSeek-R1 and GPT-4o, while Refu.($\mathcal{A}$) and Refu.($\mathcal{U}$) also marginally improve. The observation implies that reliable prompts can alleviate the unreliability issue on unsolvable cases without sacrificing the performance on solvable data. Moreover, the sequence lengths after using reliable prompts also decrease noticeably, suggesting that reliable prompts can also mitigate the overthinking issue.
c. For larger-size LLMs, the reliability of reasoning LLMs generally outperforms instruction LLMs when using reliable prompts. Conversely, smaller reasoning LLMs demonstrate inferior reliability to instruction LLMs. As in Table 3, DeepSeek-R1 and o3-mini present greater reliability compared to DeepSeek-V3 and GPT-4o in Succ.($\mathcal{U}$). However, Distill 7B and 1.5B demonstrate slightly weaker reliability than their counterparts, Qwen2.5 7B and 1.5B, particularly in Refu.($\mathcal{U}$). Such relatively small reasoning LLMs are prone to excessive invalid overthinking, with averaged generation lengths over 10k, severely undermining LLM reliability. In addition, for LLMs from the same family, larger LLMs exhibit superior reliability than small LLMs (Distill 32B vs. 14B; Distill 7B vs. 1.5B; Qwen2.5 7B vs. 1.5B) for both reasoning and instruction LLMs.
d. Despite performing well on solvable problems, relatively small LLMs can hardly detect unsolvability and show no improvement even when using reliable prompts. Whether small-sized reasoning LLMs (Distill 7B and 1.5B) or the instruction LLMs (Qwen2.5 7B and 1.5B) achieve Succ.($\mathcal{A}$) of $40\sim 60$ on solvable problems, but Succ.($\mathcal{U}$) are always almost 0 on unsolvable problems even using reliable prompts. This is because small LLMs after training overfit to generate certain answers to any mathematical problems even for unsolvable ones, thereby weakening their general instruction-following ability and preventing them from recognizing unsolvable cases.
e. For mathematical problems at different levels, LLMs exhibit high sensitivity on solvable problems but remain invariance on unsolvable ones. As in Fig. 4, for solvable problems (slv), LLMs achieve higher Succ.($\mathcal{A}$) on relatively simple high school–level MATH problems compared with the more challenging Olympiad-level AIME and AMC, and the college-level Minerva, employing either standard or reliable prompts. In contrast, for unsolvable problems (usl), Succ.($\mathcal{U}$) remain comparable across different levels, suggesting that the ability of LLMs to identify an unsolvable problem is independent of the difficulty level of the corresponding original solvable problem.
Figure 5: Illustrations of Succ.($\mathcal{U}$) of unsolvable problems regarding (a) two rewriting schemes (removal & contradiction), and (b) two difficulty levels labeled by experts (0: simple & 1: hard) on several typically used LLMs using reliable prompts. Completed results on all LLMs are presented in Supplement C.3.
4.3 Dataset Analysis
It is observed that LLMs are prone to fail on removal questions or those with difficulty level dd=1 of unsolvable cases, we thus make some analysis of our constructed unsolvable data regarding different rewriting schemes and difficulty levels as in Fig. 5.
Rewriting Schemes (Removal & Contradiction)
We present two unsolvable rewriting schemes of removal and contradiction in Sec. 3.2. As in Fig. 5 (a), we showcase the results of Succ.($\mathcal{U}$) by on two rewriting schemes on several LLMs using reliable prompts. Generally, Succ.($\mathcal{U}$) of contradiction problems is larger than removal problems, indicating that LLMs are more adept at identifying the unsolvability of contradiction cases. As we observed in generated outputs, LLMs tend to fabricate missing conditions for removal problems, resulting in verbose yet hallucinated responses, thus yielding fewer successful responses.
Difficulty Level (0 & 1)
As in Step 3 of Sec. 3.3, human experts label problems that are intuitively judged as unsolvable with a difficulty level of 0, whereas problems with a difficulty level of 1 require step-by-step reasoning to determine the unsolvability. In Fig. 5 (b), we demonstrate Succ.($\mathcal{U}$) with two difficulty levels (0: simple & 1: hard). Succ.($\mathcal{U}$) with difficulty level $d=1$ is lower than problems of $d=0$ across all LLMs, suggesting that LLMs are consistently correlated with human experts in identifying unsolvability of problems.
5 Reliability Improvements
With the preceding analysis, since such relatively small LLMs exhibit worse reliability, we further propose an alignment strategy in Sec. 5.1 to critically identify unsolvable problems, and conduct experiments to improve LLM reliability on both math and knowledge QA tasks in Sec. 5.3 and 5.2.
5.1 Alignment Strategy
Figure 6: Illustrations of (a) reliability alignment training data generation for both solvable and unsolvable problems, (b) SFT-based alignment using the obtained training data, and (c) DPO-based alignment on the obtained SFT model.
Methods In-Domain Out-of-Domain
Succ.($\mathcal{A}$) Refu.($\mathcal{A}$) Succ.($\mathcal{U}$) Refu.($\mathcal{U}$) Succ.($\mathcal{A}$) Refu.($\mathcal{A}$) Succ.($\mathcal{U}$) Refu.($\mathcal{U}$)
$s_{p}$ $s_{o}$ $s_{p}$ $s_{o}$ $s_{p}$ $s_{o}$
DeepSeek-R1
reliable 73.17 73.49 0.00 70.79 54.89 1.76 48.45 8.16 71.66 8.79 57.92
Qwen2.5-1.5B
standard 41.22 45.05 0.00 1.89 0.00 0.00 4.74 0.07 17.08 17.76 0.12
reliable 39.63 42.18 0.00 2.17 1.44 0.00 3.69 2.50 23.33 4.64 5.45
alignment-sft 48.24 48.89 9.58 17.15 22.69 9.56 3.16 10.87 31.36 25.23 12.48
alignment-dpo 46.00 46.64 5.11 21.87 27.22 8.31 1.57 8.37 25.44 30.26 7.88
Distill-1.5B
standard 40.26 41.86 0.00 2.74 0.00 0.00 0.13 0.00 11.70 0.06 0.00
reliable 38.98 39.62 0.00 2.18 0.00 0.00 0.26 0.07 11.88 0.12 0.00
alignment-sft 55.27 56.23 7.99 26.13 25.05 9.25 2.51 9.86 16.74 21.65 10.14
alignment-dpo 53.05 53.99 4.79 27.95 29.31 8.60 0.43 6.88 22.14 24.45 7.12
Table 5: Results of Success Rate (Succ.) and Refusal Rate (Refu.) on both in-domain math task and out-of-domain knowledge QA task including both solvable and unsolvable problems. Two relatively small LLMs of Distill-1.5B and Qwen2.5-1.5B are employed using standard and reliable prompts, as well as our proposed alignment strategy including both SFT and DPO. DeepSeek-R1 is the teacher model. $\dagger$ denotes that significant improvements are obtained using the alignment method over both reliable and standard baseline prompting methods.
To enhance the reliability of small LLMs $\mathcal{M}_{S}$, we first obtain alignment training data. Given the training set $\mathcal{Q}_{a}$} with open-source solvable math problems, we synthesize unsolvable problems in $\mathcal{Q}_{u}$} using the unsolvable data construction process in Sec. 3.3.333Due to the expensive cost of human checks on substantial training data, Step 3 is omitted. As demonstrated in Table 2, the pass rate of Step 3 human check is approximately 80.14%; therefore, the quality of synthesized unsolvable problems after Step 2 is regarded as applicable enough. We then generate successful and refused responses for problems in $\mathcal{Q}_{a}$} and $\mathcal{Q}_{u}$}.
Due to the limited ability of $\mathcal{M}_{S}$ to generate successful responses including high-quality reasoning paths of solvable problems or identify unsolvability of solvable problems, we employ an advanced teacher LLM $\mathcal{M}_{T}$ (DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability)) for distillation (xu2024survey). As in Fig. 6 (a), we employ the standard prompt $p_{s}$ for $\mathcal{Q}_{a}$} and the reliable prompt $p_{r}$ for $\mathcal{Q}_{u}$} to distill successful responses $\mathcal{S}$ from $\mathcal{M}_{T}$, with a sampling number $K=4$ to ensure that most problems yield at least one $\mathcal{S}$ response. To prevent LLM from refusing to answer questions it can solve, for $\mathcal{Q}_{a}$}, we first infer on small LLMs to select questions that are incorrectly answered and add to the unknown problem set $\mathcal{Q}_{n}$}. Then we employ rejection sampling (yuan2023scalingrelationshiplearningmathematical) on $\mathcal{M}_{S}$ to sample model-specific refusal responses $\mathcal{R}$ for $\mathcal{Q}_{n}$} and $\mathcal{Q}_{u}$} with $K=1$ and a 2-shot refusal prompt $p_{f}$ to accelerate sampling. Since $\mathcal{S}$ is more preferable than $\mathcal{F}$, we only need a few refusal responses to enable $\mathcal{M}_{S}$ to appropriately learn the refusal capability. The generated data are incorporated into the training set ${\mathcal{D}}^{(t)}$ to align with the reliability for $\mathcal{M}_{s}$ as follows. Prompt templates are in Supplement 15.
In Fig. 6 (b), the constructed reliable supervision data ${\mathcal{D}}^{(t)}$ is used for Supervised Fine-Tuning (SFT) to train the small LLM $\mathcal{M}_{s}$ to critically identify the unsolvability or refuse to answer while keeping their ability on solvable problems. Following the initial SFT phase, as in Fig. 6 (c), we sample and synthesize high-quality preference data based on the model’s outputs, which is subsequently utilized to further refine the model using the Direct Preference Optimization (DPO) (rafailov2024directpreferenceoptimizationlanguage) method, progressively aligning its behavior with the desired reliability preferences.
5.2 Alignment Setup
We conduct experiments on unreliable LLMs of Qwen2.5-1.5B and Distill-1.5B for reasoning reliability enhancements. We incorporate solvable problems from open-source AIME, AMC, and MATH training data into $\mathcal{Q}_{a}$}. To avoid potential data leakage due to the fundamental relation between train and test data which could be a confounder, questions in the $\mathcal{Q}_{a}$} with high similarity to the solvable test set $\mathcal{D}_{a}$ were filtered out as follows: For each $\boldsymbol{x}_{t}\in{\mathcal{Q}_{a}}$, we calculate the cosine similarity scores $\{s_{i}\}_{i=1}^{N}$ of $\boldsymbol{x}_{t}$ with all $\{\boldsymbol{x}_{i}|\boldsymbol{x}_{i}\in\mathcal{D}_{a}\}_{i=1}^{N}$, and if $\max\{s_{i}\}_{i=1}^{N}>0.7$, $\boldsymbol{x}_{t}$ is removed. The threshold $0.7$ is set empirically which filters out the top 10% problems with high similar scores.
When constructing unsolvable problems of $\mathcal{Q}_{u}$}, only one condition is extracted to maintain the size of $\mathcal{Q}_{u}$} comparable to $\mathcal{Q}_{a}$}. Distilled responses with reasoning steps are used to train the reasoning LLM Distill-1.5B, and we remove the reasoning steps to train the instruction LLM Qwen2.5-1.5B. The responses and problems are formatted and constituted to the alignment training set ${\mathcal{D}}^{(t)}$ ($|\mathcal{D}^{(t)}|=\text{10k}$). Statistics of training problems, the numbers of sampled $\mathcal{S}$ and $\mathcal{R}$ responses of $\mathcal{Q}_{a}$} and $\mathcal{Q}_{u}$}, and training details are all presented in Supplement 13.
LLMs after alignment are tested on both in-domain (ID) math tasks of ReliableMath dataset and out-of-domain (OOD) knowledge QA task derived from the KUQ dataset (amayuelas-etal-2024-knowledge). KUQ contains answerable knowledge QAs from open-source QA datasets like TriviaQA (joshi-etal-2017-triviaqa), as well as real-world unknown questions collected from crowd-source workers. Specifically, we omit categories of controversial and ambigious questions as they are not actually answerable but have various candidate answers. When inference, we also employ standard and reliable knowledge QA prompts for inference as in Supplement 15. Unlike math tasks, known knowledge questions typically do not involve explicit reasoning and can be directly evaluated using the F1 score on the outputs. Nonetheless, for unanswerable knowledge questions, it remains essential to assess whether LLMs’ reasoning process can identify the unsolvability aspects.
5.3 Alignment Experimental Results
Reliability improvements suggest that our proposed alignment methods enable LLMs not only to simply identify the unsolvability of math problems, but also to critically think and identify their knowledge or capability boundary before answering any questions. As in Table 5, both of our alignment strategy using SFT and DPO can effectively improve the reliability of two LLMs on either ID mathematical or OOD real-world knowledge QA tasks, with significant improvements of $20\sim 30$ in Succ.($\mathcal{U}$) over two baseline prompting methods. Advancements in Refu.($\mathcal{A}$) and Refu.($\mathcal{U}$) indicate that aligned LLMs acquire the refusal ability on both solvable and unsolvable/unknown problems, which is rare when using prompting methods. In addition, although LLMs using DPO alignment exhibit larger gains on Succ.($\mathcal{U}$), they present inferior generalization on other reliability metrics compared with using SFT alignment.
6 Related Work
6.1 LLM Reliability
Reliability is a foundational requirement for LLM requiring generating factually correct, informative, and consistent responses to users (liu2023trustworthy; li2024surveyhonestylargelanguage; NEURIPS2024_0d99a8c0). A unified perspective of LLM reliability is that LLMs should be able to identify questions that fall outside their scopes or are unanswerable (yin-etal-2023-large; amayuelas-etal-2024-knowledge; wang-etal-2024-enhancing). Prior work enhances the reliability of knowledge tasks by aligning LLMs with their uncertainty to confidently answer known questions and refuse unknown questions (zhang-etal-2024-r) to mitigate hallucinations (xue2024ualignleveraginguncertaintyestimations; yang2024alignment; xu2024rejection; lin2024flamefactualityawarealignmentlarge; zheng2025enhancingllmreliabilityexplicit; xue2025mlingconfcomprehensivestudymultilingual). However, current research primarily focuses on knowledge tasks. There is a void of benchmark on LLM reliability on reasoning tasks.
6.2 Mathematical Reasoning on LLMs
Math tasks have long been regarded as effective proxies of reasoning capabilities of LLMs (koncel2016mawps; miao2021diverse; tang2024mathscalescalinginstructiontuning). These tasks require LLMs to comprehend the semantics and symbols in the question, engage in problem-solving processes step by step, and present the final answers. (shao2024deepseekmathpushinglimitsmathematical; yang2024qwen25mathtechnicalreportmathematical). Open benchmarks range from school-level GSM8K (cobbe2021training) and MATH (hendrycks2021measuring) to more challenging, Olympiad- or college-level AIME (aime_1983_2024) and CollegeMath (tang2024mathscalescalinginstructiontuning). Recent advanced LLMs like DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability), emulating the slow and deliberate deep Chain-of-Thought (CoT) (NEURIPS2022_9d560961) reasoning steps (System 2 thinking), have exhibited remarkable math reasoning abilities (chen2025reasoningerasurveylong) over previous foundation and instruction LLMs like DeepSeek-V3 (deepseekai2024deepseekv3technicalreport) which execute fast and heuristic-driven generation (System 1 thinking) (li202512surveyreasoning; loo2021system12). LLMs endowed with deep reasoning capabilities for solving complex math problems are commonly referred to as reasoning LLMs (chen2025reasoningerasurveylong). However, prior work has not considered the reasoning reliability evaluation of LLMs where unreliable LLMs fabricate solutions to unsolvable/unknown/unanswerable problems, thus leading to hallucination in reasoning steps.
7 Conclusion
This work introduces a ReliableMath benchmark to systematically investigate LLM reliability on reasoning tasks. The benchmark formulates the reliability evaluation on mathematical reasoning tasks, as well as develops a dataset comprising open-source solvable math problems and high-quality unsolvable problems synthesized from our proposed data construction workflow including expert check. Experiments conducted on a series of reasoning and instruction LLMs derive several key findings and analyses. Finally, we propose an alignment strategy which can enable LLMs to critically identify unsolvable cases and significantly enhance the reliability. We believe this study can serve as a strong benchmark to inspire more valuable research to develop more reliable LLMs for future work.
8 Definition of Notations
Notation Description
Sec. 2: Reliability Evaluation
$\mathcal{M}$ LLMs.
$\mathcal{A}$ Solvable mathematical problems.
$\mathcal{U}$ Unolvable mathematical problems.
$\mathcal{S}$ Successful responses for reliability evaluation.
$\mathcal{R}$ Refused responses for reliability evaluation.
$\mathcal{S}$ Failed responses for reliability evaluation.
Sec. 3: Dataset Construction
$\mathcal{M}_{I}$ Instruction LLMs.
$\mathcal{M}_{R}$ LLMs Reasoning LLMs.
$\mathcal{D}_{r}$ ReliableMath dataset.
$\mathcal{D}_{a}$ ReliableMath solvable subset.
$\mathcal{D}_{u}$ ReliableMath unsolvable subset.
$\boldsymbol{x}_{i}$ The $i$-th question sample from $\mathcal{D}_{a}$.
$\boldsymbol{\hat{y}}_{i}$ The ground-truth answer to $\boldsymbol{x}_{i}$ in $\mathcal{D}_{a}$.
$\boldsymbol{y}_{i}$ The LLM generated response to $\boldsymbol{x}_{i}$.
$\boldsymbol{c}_{i}^{j}$} The $j$-th extracted condition of $\boldsymbol{x}_{i}$.
$\boldsymbol{\tilde{x}}_{i}^{j}$} The $j$-th rewritten question of $\boldsymbol{x}_{i}$.
$\boldsymbol{\tilde{c}}_{i}^{j}$} The rewritten condition by model verification of $\boldsymbol{\tilde{x}}_{i}^{j}$} where $\boldsymbol{\tilde{c}}_{i}^{j}$} entails $\boldsymbol{c}_{i}^{j}$}.
$\boldsymbol{\tilde{a}}_{i}^{j}$} The unsolvable reason analysis by model verification of $\boldsymbol{\tilde{x}}_{i}^{j}$}.
${d}_{i}^{j}$} The difficulty level by human evaluation of $\boldsymbol{\tilde{x}}_{i}^{j}$}.
Sec. 5: Reliability Improvement
$\mathcal{M}_{T}$ Advanced teacher LLM to be distilled.
$\mathcal{M}_{S}$ Relatively small LLMs to be improved.
${\mathcal{Q}_{a}}^{(t)}$ Solvable problem training set.
${\mathcal{Q}_{u}}^{(t)}$ Unsolvable problem training set.
${\mathcal{D}}^{(t)}$ Alignment training set with problems and sampled responses.
$p_{s}$ Standard CoT-based mathematical problem-solving prompt.
$p_{r}$ Mathematical problem-solving prompt with refusal instruction.
$p_{u}$ Mathematical problem-solving prompt with instruction to indicate unsolvability.
Table 6: Summarized notations in this work.
The definitions of the notations in this work are summarized in Table 6.
9 Dataset Details
9.1 Solvable Data Collection
  • AIME 2024 (aimo2024aime)  [^4] https://huggingface.co/datasets/AI-MO/aimo-validation-aime serves as a benchmark comprising exceptionally challenging mathematical problems from the American Invitational Mathematics Examination (AIME). These problems are specifically designed to evaluate advanced problem-solving abilities and are considerably more difficult than standard high school mathematics questions. Solving these items requires sophisticated reasoning and the application of advanced strategies.
  • AMC dataset (aimo2024amc)  [^5] https://huggingface.co/datasets/AI-MO/aimo-validation-amc serves as an internal validation set during participation in the AIMO progress prize competition. The dataset contains 83 problems from AMC12 2022 and AMC12 2023 extracted from the AOPS wiki page  [^6] https://artofproblemsolving.com/wiki/index.php/ AMC_12_Problems_and_Solutions.
  • MinervaMath (lewkowycz2022solving)  [^7] https://huggingface.co/datasets/math-ai/minervamath includes 272 college-level problems in physical, engineering, chemistry, economics, and other sciences that require quantitative reasoning.
  • MATH (hendrycks2021measuring)  [^8] https://github.com/hendrycks/math/ is a challenging mathematical dataset with competitive mathematics problems, consisting of 7,500 training samples and 5,000 test samples. Each problem in MATH also has a full step-by-step solution which can be used to teach models to generate answer derivations and explanations across several subjects, including algebra, geometry, number theory, counting and probability, calculus, etc.
9.2 Data Format
The data format of solvable subset is demonstrated in Fig. 7.
$\{$ ‘‘id’’: // data index
‘‘data_source’’: // data source
‘‘question’’: // original question
‘‘ground_truth’’: // original ground truth
Figure 7: Solvable data format in ReliableMath dataset.
Figure 8: Workflow of ReliableMath unsolvable dataset construction.
10 Ethics Statement
In this paper, we introduce a ReliableMath dataset derived from the publicly available dataset. Most human annotators are experts with at least a bachelor’s degree in computer science or math, who are from crowd-sourcing platforms. We meticulously adhered to legal and ethical standards throughout the data collection process, prioritizing privacy and obtaining informed consent. We have constructed ReliableMath dataset based on existing mathematical test datasets, with the foundational data not involving any knowledge that poses safety hazards, privacy violations, or ethical breaches. Consequently, the ReliableMath dataset we have built does not contain any harmful information. Experts were furnished with comprehensive details regarding the study’s objectives, data collection methodologies, and associated risks or benefits. They are paid a wage higher than the local average hourly rate, which is provided by the employing partner. They were afforded the opportunity to seek clarification and voluntarily provide consent before their involvement. All collected data were exclusively utilized for research purposes.
11 Potential Social Impact
This work can serve as a strong benchmark to advance AI applications in the field of mathematical reasoning, enhancing the reliability of AI applications and reducing hallucinations, with great positive social impacts. Our benchmark can also strengthen LLMs’ reliability in some educational applications for the community. However, since we introduce an unsolvable mathematical dataset, if these problems are misused to query LLMs, they could lead to excessively long outputs, causing the LLMs to crash and potentially resulting in attacks on the systems.
12 Human Evaluation Instruction
Figure 9: Examples of human evaluated data.
12.1 Description of Task
The unsolvable dataset for human evaluation is based on high-quality and original math problems from test datasets, including AIME, AMC, Math, and Minerva. The problems are rewritten to be unsolvable through two methods: a. removing key mathematical conditions, and b. adding conditions that contradict the given information.
Then, the original question, ground-truth answer, and the auxiliary information, including the rewriting process and the modified positions, will be presented to the evaluator to assist in the manual annotation process of verifying whether the question is unsolvable, thereby improving the quality of human evaluation. The detailed format of data presented to the human annotator will be explained in Section 12.2.
The evaluator’s goal is to combine the above-mentioned information to determine whether the constructed unsolvable problem is indeed unsolvable. The detailed evaluation principles will be explained in Section 12.3.
12.2 Evaluation Data Format and Explanation
As shown in Fig. 10, we illustrate the format of annotated data and the explanation of items. The "data_source", "question", and "ground_truth" are from the original dataset, Human annotators are required to label two fields: "human_eval" and "difficulty_level".
$\{$ ‘‘unsol_id’’: // unsolvable data index
‘‘data_source’’: // data source: AIME, AMC, Math ot Minerva
‘‘question’’: // original question
‘‘ground_truth’’: // original ground truth
‘‘rewritten_condition’’: // rewritten conditions
‘‘rewritten_question’’: // rewritten question after removing conditions or adding contradictory ones
‘‘unsolvable_reason’’: // unsolvable reason analysis of rewritten question
‘‘human_check’’: // human check result. Determine whether the rewriting problem is unsolvable, unsolvable: 1, solvable: 0
‘‘difficulty_eval’’: // Determining the difficulty of whether this problem is unsolvable
Figure 10: Illustration of human annotation results.
12.3 Principle of Evaluation
In this section, we provide a detailed explanation of the specific evaluation criteria for human annotators to label the data’s ‘human_eval‘ and ‘difficulty_level‘ fields.
human_check
Given the example data, the evaluator must first determine whether the "rewritten_question" meets the following requirements, which includes three points:
  • If too many prerequisites of original questions are removed or unreasonable conditions are added, even leading to a problem that does not meet the definition of a mathematical problem, then the score is 0. As exemplified in Fig. 9 a.
  • Combining "rewritten_condition", which illustrates the altered part, determines whether the rewritten question "rewritten_question" has changed the conditions of the original question "question". If the rewritten question does not meet the rewritten requirements (e.g., without a condition removed or added), then the score is 0.
  • Combining "unsolvable_reason", the original question ("question"), the answer ("ground_truth"), determine whether the rewritten question is indeed unsolvable. If it is still solvable, then the score is 0; otherwise, it is 1.
    To reduce the difficulty of evaluation, the evaluator will be not required to solve the problem, they only need to combine the "unsolvable_reason" and the original question solution to judge.
    For example, if some questions require the maximum value, the rewritten answer is positive infinity, still solvable, but if the question is the maximum real number solution, then it is unsolvable, as exemplified in Fig. 9 b.
difficulty_eval
Determine the difficulty of judging whether this question is unsolvable. If the output of "human_check" is 1, then further evaluate the difficulty, and the evaluator can combine "unsolvable_reason" to judge.
  • If the unsolvable reason of the rewritten question is relatively simple and shallow, the difficulty is set to 0, as exemplified in Fig. 9 c.
    For example, (1) the problem having a $log(x)$, with the added contradiction condition $x=0$; (2) the geometric graph of the questions does not match, which is relatively easy to see that the question is definitely unsolvable.
  • If it requires a certain level of reasoning to determine that the question is unsolvable, then the difficulty is set to 1.
    For example, the original question has been removed of a constraint condition, and it requires reasoning (e.g., cost the annotator several minutes of thinking) to determine that it is unsolvable.
Dataset Ori. $\mathcal{D}_{a}$ Step 1 Step 2 Step 3 ($\mathcal{D}_{u}$)
Rmv. Cod. Total Rmv. Cod. Total Rmv. Cod. Total
AIME 30 30 85 85 170 70 71 141 67 65 132
AMC 83 83 241 241 482 185 192 377 131 164 295
MATH500 500 100 273 273 546 233 216 449 154 164 318
Minerva 272 100 254 254 508 207 201 408 185 172 357
Total 885 313 853 853 1706 695 680 1375 537 565 1102
Table 7: Data statistics after each step. Rmv. and Cod. denote Removal and Contradiction respectively. Ori. is the number of original datasets and $\mathcal{D}_{a}$ denotes the number of our employed solvable problems.
Dataset Step 3 ($\mathcal{D}_{u}$) (diff. 0 / diff. 1)
Rmv. Cod. Total
AIME 67 (24/43) 65 (22/43) 132 (46/86)
AMC 131 (47/84) 164 (60/104) 295 (107/188)
MATH500 154 (77/77) 164 (75/89) 318 (152/166)
Minerva 185 (87/96) 172 (76/96) 357 (163/192)
Total 537 (235/300) 565 (233/332) 1102 (468/632)
Table 8: Data statistics for Step 3 where we present the number of two difficulty levels (0/1).
13 Experiments
13.1 Model Details
We attach the model links of employed LLMs as follows:
13.2 Completed Experimental Results
The completed experiments on respective ReliableMath subsets of AIME, AMC, MATH, and Minerva are presented in Table 11, 12, 13, 14, and Fig. 12.
13.3 Training Settings
All the LLMs are trained using Llama-Factory (zheng2024llamafactory).  [^17] https://github.com/hiyouga/LLaMA-Factory During training, ADAM parameter update is used in a mini-batch mode with batch size = 16. The initial learning rate of 1e-5 is utilized with the 0.05 warm-up ratio and 0.01 weight decay of the ADAM optimizer. When training the models, we set epochs to 3 to ensure that all the models can be trained to convergence. All LLMs are quantified using brain floating point (bf16) to load and save parameters. The vLLM library (kwon2023efficient)  [^18] https://github.com/vllm-project/vllm is utilized to accelerate the generation.
Dataset $\mathcal{Q}_{a}$} Step 1 Step 2 $\mathcal{Q}_{u}$}
AIME 975 1950 961
AMC 3264 6528 3185
MATH 2298 4596 961
Total 6537 13074 5085
Table 9: Training data statistics after each step. We only employ the data in MATH whose difficulty level is 5.
Succussful Refused
Solvable 5344 481
Unsolvable 4190 787
Table 10: Numbers of sampled successful and refused responses for solvable and unsolvable problems in ${\mathcal{D}}^{(t)}$.
Figure 11: Example of the generation of a removal question on DeepSeek-R1. Although LLM can perceive that the question seems missing some conditions but still attempt to fabricate reasoning steps by assume some conditions which are highlighted.
Figure 12: Results of Success Rate (Succ.) on different test sets (AIME, AMC, MATH, Minerva) on both solvable (slv) and unsolvable (usl) subsets using standard (std) and reliable (real) prompts, respectively.
LLMs Prompt AIME Solvable ($\mathcal{A}$) AIME Unsolvable ($\mathcal{U}$) Succ. Refu.
Succ. ($\mathcal{A}$) Refu.($\mathcal{A}$) Len. Succ.($\mathcal{U}$) Refu.($\mathcal{U}$) Len.
$s_{p}$ $s_{o}$ $s_{p}$ $s_{o}$
DeepSeek-R1 standard 83.33 83.33 0.00 7.59k 59.85 0.00 0.00 11.07k 56.62 0.00
reliable 70.00 70.00 0.00 7.57k 69.70 39.39 3.00 8.53k 62.27 1.50
o3-mini standard 60.00 63.33 0.00 3.46k 16.67 0.00 0.00 8.47k 35.00 0.00
reliable 73.33 73.33 0.00 3.89k 35.61 25.00 5.00 5.21k 51.81 2.50
Distill-32B standard 63.33 66.67 0.00 8.34k 19.70 0.00 0.00 20.63k 37.42 0.00
reliable 63.33 63.33 0.00 12.57k 37.88 25.76 2.00 15.68k 47.58 1.00
Distill-14B standard 60.00 60.00 0.00 12.83k 26.52 0.00 0.00 26.30k 36.63 0.00
reliable 50.00 53.33 0.00 15.46k 51.52 34.85 1.00 20.26k 47.42 0.50
Distill-7B standard 43.33 46.67 0.00 11.52k 2.27 0.00 0.00 11.13k 23.07 0.00
reliable 36.67 36.67 0.00 12.00k 3.79 0.00 0.00 11.80k 19.28 0.00
Distill-1.5B standard 23.33 26.67 0.00 13.56k 2.27 0.00 0.00 13.36k 13.07 0.00
reliable 20.00 20.00 0.00 13.88k 1.52 0.00 0.00 13.54k 10.38 0.00
DeepSeek-V3 standard 46.67 46.67 0.00 3.10k 22.73 0.00 0.00 3.51k 29.02 0.00
reliable 46.67 46.67 0.00 3.01k 37.12 18.18 0.00 3.01k 37.16 0.00
GPT-4o standard 0.00 6.67 0.00 0.77k 6.06 0.00 0.00 0.74k 3.18 0.00
reliable 10.00 10.00 0.00 0.69k 21.21 16.67 5.00 0.93k 14.47 2.50
Qwen2.5-7B standard 23.33 30.00 0.00 1.28k 3.03 0.00 0.00 1.23k 14.10 0.00
reliable 13.33 16.67 0.00 1.49k 0.76 0.76 0.00 1.46k 7.88 0.00
Qwen2.5-1.5B standard 6.67 6.67 0.00 0.89k 2.27 0.00 0.00 0.95k 3.90 0.00
reliable 10.00 10.00 0.00 0.95k 6.06 0.00 0.00 0.89k 6.51 0.00
Table 11: Reliability evaluations on ReliableMath AIME Solvable and Unsolvable subsets.
LLMs Prompt AMC Solvable ($\mathcal{A}$) AMC Unsolvable ($\mathcal{U}$) Succ. Refu.
Succ. ($\mathcal{A}$) Refu.($\mathcal{A}$) Len. Succ.($\mathcal{U}$) Refu.($\mathcal{U}$) Len.
$s_{p}$ $s_{o}$ $s_{p}$ $s_{o}$
DeepSeek-R1 standard 91.57 92.77 0.00 5.23k 68.14 0.00 0.00 7.92k 63.12 0.00
reliable 89.16 90.36 0.00 4.90k 70.51 51.53 2.00 5.34k 75.39 1.00
o3-mini standard 84.34 85.54 0.00 2.22k 17.29 0.34 0.00 5.56k 46.88 0.00
reliable 86.75 87.95 0.00 2.90k 28.47 30.51 1.00 5.67k 58.42 0.50
Distill-32B standard 84.34 85.54 0.00 5.81k 33.56 0.00 0.00 17.69k 50.86 0.00
reliable 80.72 81.93 0.00 5.65k 50.85 36.27 0.00 13.14k 62.45 0.00
Distill-14B standard 84.34 85.54 0.00 7.38k 41.69 0.00 0.00 20.43k 52.89 0.00
reliable 79.52 80.72 0.00 7.74k 61.69 45.42 0.00 14.52k 66.84 0.00
Distill-7B standard 80.72 80.72 0.00 6.92k 0.68 0.00 0.00 7.08k 40.53 0.00
reliable 63.86 63.86 0.00 7.92k 2.04 0.00 0.00 7.99k 32.44 0.00
Distill-1.5B standard 45.78 45.78 0.00 10.18k 2.38 0.00 0.00 10.14k 23.48 0.00
reliable 45.78 45.78 0.00 10.12k 1.70 0.00 0.00 10.04k 23.32 0.00
DeepSeek-V3 standard 78.31 80.72 0.00 2.17k 36.95 0.00 0.00 2.58k 48.99 0.00
reliable 81.93 83.13 0.00 2.00k 50.17 29.15 2.00 2.19k 61.09 1.00
GPT-4o standard 40.96 45.78 0.00 1.15k 10.85 0.00 0.00 1.01k 24.39 0.00
reliable 36.14 44.58 0.00 0.74k 29.49 30.17 10.00 0.90k 35.09 5.00
Qwen2.5-7B standard 40.96 49.40 0.00 1.08k 2.38 0.00 0.00 1.06k 23.18 0.00
reliable 45.78 57.83 0.00 0.92k 5.76 1.69 0.00 0.94k 27.77 0.00
Qwen2.5-1.5B standard 45.78 54.22 0.00 0.82k 2.03 0.00 0.00 0.84k 25.51 0.00
reliable 39.76 46.99 0.00 0.83k 1.36 0.00 0.00 0.85k 22.02 0.00
Table 12: Reliability evaluations on ReliableMath AMC Solvable and Unsolvable subsets.
LLMs Prompt MATH Solvable ($\mathcal{A}$) MATH Unsolvable ($\mathcal{U}$) Succ. Refu.
Succ. ($\mathcal{A}$) Refu.($\mathcal{A}$) Len. Succ.($\mathcal{U}$) Refu.($\mathcal{U}$) Len.
$s_{p}$ $s_{o}$ $s_{p}$ $s_{o}$
DeepSeek-R1 standard 91.00 93.00 0.00 3.12k 58.49 0.00 0.00 5.76k 60.62 0.00
reliable 93.00 93.00 0.00 2.71k 71.38 59.43 3.00 3.79k 79.20 1.50
o3-mini standard 86.00 89.00 0.00 1.25k 13.52 0.00 0.00 6.18k 47.13 0.00
reliable 87.00 91.00 0.00 0.82k 23.27 32.70 0.00 5.21k 58.49 0.00
Distill-32B standard 88.00 90.00 0.00 4.31k 29.87 0.00 0.00 12.13k 51.97 0.00
reliable 85.00 87.00 0.00 4.52k 55.35 49.37 0.00 7.81k 69.18 0.00
Distill-14B standard 83.00 84.00 0.00 4.83k 38.68 0.00 0.00 14.47k 51.42 0.00
reliable 82.00 82.00 0.00 5.15k 60.69 52.20 0.00 9.35k 69.22 0.00
Distill-7B standard 84.00 85.00 0.00 4.08k 2.52 0.00 0.00 4.54k 42.88 0.00
reliable 83.00 84.00 0.00 4.43k 1.89 0.00 0.00 4.89k 42.22 0.00
Distill-1.5B standard 65.00 67.00 0.00 6.50k 4.40 0.00 0.00 7.43k 34.10 0.00
reliable 63.00 64.00 0.00 7.14k 3.77 0.00 0.00 7.77k 32.70 0.00
DeepSeek-V3 standard 88.00 89.00 0.00 1.15k 25.79 0.00 0.00 1.80k 50.70 0.00
reliable 88.00 90.00 0.00 1.16k 42.14 42.45 0.00 1.41k 65.65 0.00
GPT-4o standard 68.00 72.00 0.00 0.55k 9.75 0.00 0.00 0.74k 37.44 0.00
reliable 66.00 70.00 1.00 0.51k 29.56 37.74 6.00 0.49k 50.83 3.50
Qwen2.5-7B standard 75.00 79.00 0.00 0.67k 0.63 0.00 0.00 0.70k 38.66 0.00
reliable 77.00 80.00 0.00 0.69k 0.63 0.63 0.00 0.77k 39.56 0.00
Qwen2.5-1.5B standard 65.00 68.00 0.00 0.58k 0.63 0.00 0.00 0.65k 33.41 0.00
reliable 67.00 68.00 0.00 0.59k 2.83 0.63 0.00 0.61k 34.62 0.00
Table 13: Reliability evaluations on ReliableMath MATH Solvable and Unsolvable subsets.
LLMs Prompt Minerva Solvable ($\mathcal{A}$) Minerva Unsolvable ($\mathcal{U}$) Succ. Refu.
Succ. ($\mathcal{A}$) Refu.($\mathcal{A}$) Len. Succ.($\mathcal{U}$) Refu.($\mathcal{U}$) Len.
$s_{p}$ $s_{o}$ $s_{p}$ $s_{o}$
DeepSeek-R1 standard 39.00 39.00 0.00 3.12k 54.06 0.00 0.00 4.27k 33.02 0.00
reliable 41.00 41.00 0.00 2.87k 70.87 59.38 0.00 2.58k 53.06 0.00
o3-mini standard 29.00 29.00 0.00 0.59k 12.61 0.00 0.00 1.90k 17.65 0.00
reliable 38.00 38.00 2.00 0.53k 22.13 26.89 0.00 1.91k 31.26 1.00
Distill-32B standard 40.00 42.00 0.00 3.97k 36.13 0.00 0.00 11.82k 29.54 0.00
reliable 40.00 40.00 0.00 2.81k 53.78 45.66 0.00 5.26k 44.86 0.00
Distill-14B standard 36.00 37.00 0.00 5.57k 35.29 0.00 0.00 12.92k 27.07 0.00
reliable 32.00 32.00 0.00 3.31k 59.94 46.50 0.00 6.03k 42.61 0.00
Distill-7B standard 25.00 25.00 0.00 6.00k 1.68 0.00 0.00 6.10k 12.92 0.00
reliable 32.00 32.00 0.00 4.94k 1.40 0.84 0.00 5.04k 16.56 0.00
Distill-1.5B standard 16.00 18.00 0.00 9.15k 1.68 0.00 0.00 9.62k 8.92 0.00
reliable 15.00 16.00 0.00 9.63k 1.40 0.00 0.00 9.73k 8.10 0.00
DeepSeek-V3 standard 34.00 35.00 0.00 0.58k 23.53 0.28 0.00 0.83k 23.20 0.00
reliable 35.00 35.00 0.00 0.49k 40.90 47.90 1.00 0.56k 39.70 0.50
GPT-4o standard 32.00 32.00 0.00 0.51k 11.76 0.28 0.00 0.66k 19.01 0.00
reliable 33.00 34.00 1.00 0.48k 25.77 38.66 6.00 0.47k 32.86 3.50
Qwen2.5-7B standard 24.00 27.00 0.00 0.65k 0.84 0.00 0.00 0.69k 12.96 0.00
reliable 21.00 25.00 0.00 0.67k 4.76 6.16 0.00 0.70k 14.23 0.00
Qwen2.5-1.5B standard 24.00 26.00 0.00 0.69k 2.80 0.00 0.00 0.72k 13.20 0.00
reliable 21.00 22.00 0.00 0.74k 0.84 3.92 0.00 0.75k 11.94 0.00
Table 14: Reliability evaluations on ReliableMath Minerva Solvable and Unsolvable subsets.
Figure 13: Illustrations of Succ.($\mathcal{U}$) of unsolvable problems regarding (a) two rewriting schemes (removal & contradiction), and (b) two difficulty levels labeled by experts (0: simple & 1: hard) on several typically used LLMs using reliable prompts.
14 Related Works
14.1 LLM Reliability
The primary function of large language models is to generate reliable responses for users, meaning that the content should be factually accurate, informative, consistent, and trustworthy to users (liu2023trustworthy; li2024surveyhonestylargelanguage; NEURIPS2024_0d99a8c0). Reliability is a foundational requirement for LLM alignment because unreliable outputs would negatively undermine almost all LLM applications and mislead users (wang2023aligning), such as decision-making (zhang2020effect) and problem-solving (garcia2016elementary) contexts.
Although the definition and categorization of LLM reliability are many-sided and depend on specific tasks, a unified perspective is that LLMs should be able to identify questions that fall outside their knowledge scope or are inherently meaningless and unanswerable (yin-etal-2023-large; amayuelas-etal-2024-knowledge; wang-etal-2024-enhancing). The most direct approach to enhancing the reliability of knowledge-based tasks is to teach LLMs to refuse to answer questions that are beyond their knowledge boundaries (zhang-etal-2024-r), more generally, express what they know and do not know, rather than providing hallucinated answers with confidence. Recent works tackle the problem by aligning LLMs with factuality and uncertainty (xue2024ualignleveraginguncertaintyestimations; yang2024alignment; xu2024rejection; lin2024flamefactualityawarealignmentlarge; zheng2025enhancingllmreliabilityexplicit; xue2025mlingconfcomprehensivestudymultilingual).
However, current research on reliability primarily focuses on knowledge-intensive tasks (yin-etal-2023-large; amayuelas-etal-2024-knowledge), while studies on reasoning tasks remain scarce. Such methods or evaluations are model-specific. For more complex reasoning tasks, users not only expect the LLMs to refuse unknown questions, but also that LLMs can determine data-specific solvability. There is a void of reliable benchmarks and datasets, widely acknowledged evaluation standards, and effective improvement methods in this field.
14.2 Reasoning LLM
Recent advanced LLMs like OpenAI-o1 (openai2024openaio1card) and DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability) have gained significant research interests in improving their reasoning capabilities (chen2025reasoningerasurveylong; openai2024o1). Such LLMs, also referred to reasoning LLMs, are designed to emulate the slow and deliberate Chain-of-Thought (CoT) NEURIPS2022_9d560961 reasoning steps (System 2 thinking), while previous foundation and instruction-tuned LLMs like GPT-4o (openai2024gpt4o), DeepSeek-V3 (deepseekai2024deepseekv3technicalreport), and Llama-3 (grattafiori2024llama3herdmodels) execute fast and heuristic-driven generation (System 1 thinking) (li202512surveyreasoning; loo2021system12). The improvements of reasoning LLMs have been primarily focused on the following three aspects: 1) Many studies have employed reinforcement learning techniques like PPO (schulman2017proximalpolicyoptimizationalgorithms) and GRPO (deepseekai2025deepseekr1incentivizingreasoningcapability) to incentivize LLMs to explore more effective reasoning strategies (shao2024deepseekmathpushinglimitsmathematical; cui2025processreinforcementimplicitrewards); 2) Researchers also endeavor to develop carefully curated, high-quality reasoning datasets to promote reasoning performance (ye2025limoreasoning; muennighoff2025s1simpletesttimescaling). 3) Test-time scaling law (snell2024scalingllmtesttimecompute) also inspires to elicit LLMs to produce long-CoT contents with various search algorithms during inference time (gandhi2024streamsearchsoslearning).
Although reasoning LLMs have showcased remarkable performance on various benchmarks, they have revealed several shortcomings. Numerous investigations (wang2025harnessingreasoningeconomysurvey; chen2025think23overthinkingo1like; qu2025surveyefficientreasoninglarge; liu2025efficientinferencelargereasoning) identified the “overthinking” issue where reasoning LLMs always generate massive and verbose thinking steps even for a simple or meaningless query, resulting in potential hallucinations and severe redundancy of tokens. Additionally, the negative impact of long-form reasoning on reliability remains underexplored and is still subject to critical skepticism.
14.3 Mathematical Tasks for LLMs
Mathematical tasks have long been regarded as effective proxies for both enhancing and assessing the reasoning capabilities of LLMs (koncel2016mawps; miao2021diverse; tang2024mathscalescalinginstructiontuning). These tasks require LLMs to comprehend the semantics and symbols in the question, engage in problem-solving processes step by step, and present the final answers. Moreover, the verifiability of the final answer facilitates straightforward evaluation and reward design for rule-based RL methods (cobbe2021training; shao2024deepseekmathpushinglimitsmathematical; yang2024qwen25mathtechnicalreportmathematical).
In previous years, GSM8K (cobbe2021training) and MATH (hendrycks2021measuring) have been utilized as prominent mathematical benchmarks for LLM training and evaluation. Recently, to further facilitate the reasoning capabilities of LLMs, researchers have paid attention to more challenging, Olympiad- or college-level mathematical problems like AIME (aime_1983_2024) and CollegeMath (tang2024mathscalescalinginstructiontuning). Therefore, extensive mathematical training corpora have been proposed (li2024numinamath; yu2023metamath; azerbayev2024llemmaopenlanguagemodel; liu2025acemathadvancingfrontiermath).
Moreover, concerns are raised that LLMs genuinely master reasoning versus pattern memorization, and many studies have sought to evaluate true mathematical reasoning abilities through various verification methods (zeng2025mrgsmk; mirzadeh2024gsmsymbolicunderstandinglimitationsmathematical; li2024gsmpluscomprehensivebenchmarkevaluating; fan2025missingpremiseexacerbatesoverthinking). Nevertheless, prior work has not considered the scenario in which LLMs attempt to generate solutions when confronted with unknown or unsolvable problems, posing potential risks of hallucination and undermining reliability. There is also a lack of corresponding datasets containing unsolvable mathematical problems for thorough assessments.
15 Prompt Templates
Template of Standard Mathematical Problem-Solving Prompt You are a math problem solver. Solve the following math problem and provide the final answer in the specified format.
[…]
### Mathematical Question ###: <math_question>
Figure 14: Template of Standard Mathematical Problem-Solving Prompt.
Template of Reliable Mathematical Problem-Solving Prompt You are a math problem solver. Solve the following math problem and provide the final answer in the specified format.
[…]
### Mathematical Question ###: <math_question>
Figure 15: Template of Reliable Mathematical Problem-Solving Prompt.
Prompt Template for Condition/Primise Extraction Instruction You are an excellent extractor.
[…]
### Mathematical Question ###: <original_math_question>
### Extracted Condition ###:
Figure 16: Prompt Template for Condition/Primise Extraction Instruction.
Prompt Template for Condition Removal Rewriting Instruction You are a good mathematical question rewriter.
Given a mathematical question ### Original Mathematical Question ###, the answer to the question is ### Original Answer ###, and ### Extracted Condition ### is a key mathematical condition extracted from the question.
Your objective is to draw inspiration from the ### Original Mathematical Question ### and ### Original Answer ### to rewrite a new math question ### Rewritten Mathematical Question ### with the following two requirements.
1) The new mathematical question ### Rewritten Mathematical Question ### need to (remove the key condition / add conditions that contradict to the condition) ### Extracted Condition ### from the ### Original Mathematical Question ###. Maintain the rest of the conditions of ### Original Mathematical Question ### unchanged.
2) After (removing the key condition / adding conditions that contradict to the condition) ### Extracted Condition ###, the new mathematical question ### Rewritten Mathematical Question ### should be unsolvable.
### Original Mathematical Question ###: <original_math_question>
### Original Answer ###: <original_answer>
### Extracted Condition ###: <extracted_condition>
Think step by step to explain how you make the problem unsolvable after (removing the condition / adding the contradicted condition). You only need to directly output your analysis after ### Analysis ###. Simplify your output as much as possible.
### Analysis ###: <analysis>
Given your analysis in ### Analysis ### about how you make the problem unsolvable after (removing the condition / adding the contradicted condition), you only need to directly output your new mathematical question after ### Rewritten Mathematical Question ###.
### Rewritten Mathematical Question ###:
Figure 17: Prompt Template for Condition Removal Rewriting Instruction.
Prompt Template for Model Verification Instruction You are an excellent verifier.
Given a mathematical question ### Original Mathematical Question ###, the answer to the question is ### Original Answer ###.
Your objective is to determine whether the ### Rewritten Mathematical Question ### (has removed some key conditions from / has added some conditions that contradict conditions of) ### Original Mathematical Question ### while keeping others unchanged.
You only need to directly output your judgement (“True” or “False”) after ### Your judgement (True or False) ###.
### Original Mathematical Question ###: <original_math_question>
### Original Answer ###: <original_answer>
### Rewritten Mathematical Question ###: <rewritten_math_question>
### Your judgement (True or False) ###:
Figure 18: Prompt Template for Model Verification Instruction.
Prompt Template for Condition Extraction Instruction You are an excellent verifier.
Given a mathematical question ### Original Mathematical Question ###, the answer to the question is ### Original Answer ###.
Your objective is to point out which key mathematical condition that ### Rewritten Mathematical Question ### (removed from / added contradicts to the conditions of) ### Original Mathematical Question ###.
You only need to directly output your response after ### Rewritten Condition ###. Simplify your output as much as possible.
### Original Mathematical Question ###: <original_math_question>
### Original Answer ###: <original_answer>
### Rewritten Mathematical Question ###: <rewritten_math_question>
### Rewritten Condition ###:
Figure 19: Prompt Template for Condition Extraction Instruction.
Prompt Template for Unsolvable Verification Instruction You are an excellent verifier.
Given a mathematical question ### Original Mathematical Question ###, the answer to the question is ### Original Answer ###.
Your objective is to determine whether the ### Rewritten Mathematical Question ### (has removed some key conditions from / has added some conditions that contradict the conditions of) ### Original Mathematical Question ###, thereby making ### Rewritten Mathematical Question ### unsolvable.
### Original Mathematical Question ###: <original_math_question>
### Original Answer ###: <original_answer>
### Rewritten Mathematical Question ###: <rewritten_math_question>
Think step by step to analyze if ### Rewritten Mathematical Question ### is unsolvable. Explain your reason and directly output after ### Unsolvable Analysis ###. Summarize and simplify your output.
### Unsolvable Analysis ###: <analysis>
Based on your analysis, you only need to directly output your judgement (“True” or “False”) after ### Your judgement (True or False) ### about if the ### Rewritten Mathematical Question ### is unsolvable.
### Your judgement (True or False) ###:
Figure 20: Prompt Template for Unsolvable Verification Instruction.
Prompt Template for Unsolvable Reason Instruction You are an excellent verifier.
Given a mathematical question ### Original Mathematical Question ###, the answer to the question is ### Original Answer ###, and ### Rewritten Mathematical Question ### (has removed some key conditions from / has added some conditions that contradict the conditions of) ### Original Mathematical Question ###, thereby making ### Rewritten Mathematical Question ### unsolvable.
### Original Mathematical Question ###: <original_math_question>
### Original Answer ###: <original_answer>
### Rewritten Mathematical Question ###: <rewritten_math_question>
### Unsolvable Analysis ###: <analysis>
Based on your analysis, you only need to directly output your reason after ### Unsolvable Reason ### about why the ### Rewritten Mathematical Question ### is unsolvable. Summarize and simplify your output.
### Unsolvable Reason ###:
Figure 21: Prompt Template for Unsolvable Reason Instruction.
Template of Model-based Process Evaluation for Solvable Problems ### Problem ###: <math_question>
### Ground-Truth Answer ###: <ground truth answer>
Given a reasoning path to the problem generated by the model.
### Reasoning Path ###: <reasoning_path>
[…]
Figure 22: Template of Model-based Process Evaluation for Solvable Problems.
Template of Model-based Process Evaluation for Unsolvable Problems (Given an unsolvable problem which misses an key condition and thus becomes unsolvable. / Given an unsolvable problem which has contradiction to the condition and thus becomes unsolvable.)
### Unsolvable Problem ###: <unsolvable_question>
### Condition ###: <ground truth answer>
Given a reasoning path to the problem generated by the model.
### Reasoning Path ###: <reasoning_path>
[…]
Figure 23: Template of Model-based Process Evaluation for Unsolvable Problems.
Template of Reliable Mathematical Problem-Solving Prompt You are a math problem solver. Solve the following math problem and provide the final answer in the specified format.
[…]
### Question ###: <few-shot example 1 question> ### Anwer ###: <few-shot example 1 refusal response> ### Mathematical Question ###: <math_question>
Figure 24: Template of Refusal Mathematical Problem-Solving Prompt.
Template of Standard Question-Answering Prompt You are an excellent question responder. Answer the question based on your knowledge.
[…]
Figure 25: Template of Standard Question-Answering Prompt.
Template of Reliable Question-Answering Prompt You are an excellent question responder. Answer the question based on your knowledge.
[…]
Figure 26: Template of Reliable Question-Answering Prompt.
Appendix A Author/Affiliation Options as set forth by MIT Press
Option 1. Author’s address is underneath each name, centered.
First Author
First Affiliation
First Address 1
First Address 2
first.email@example.com
Second Author
Second Affiliation
Second Address 1
Second Address 2
second.email@example.com
Third Author
Third Affiliation
Third Address 1
Third Address 2
third.email@example.com
Option 2. Author’s address is linked with superscript characters to its name, author names are grouped, centered.
First AuthorSecond AuthorThird Author
First Affiliation
First Address 1
First Address 2
first.email@example.com
Second Affiliation
Second Address 1
Second Address 2
second.email@example.com
Third Affiliation
Third Address 1
Third Address 2
third.email@example.com